OpenStudy (anonymous):

please help i tried answering these questions already and got them wrong !:(

OpenStudy (teddyiswatshecallsme):

What are the questions?

OpenStudy (anonymous):

What number needs to be added to complete the square of x2 – 3x = 7? Select one: a. 9/4 b. 3/2 c.-3/2 d.-9/4

OpenStudy (anonymous):

What is the vertex of y = x2 – 2x = 7? Select one: a. (1, –8) b. (–1, 8) c. (–1, –8) d. (1, 8)

OpenStudy (ybarrap):

$$ x^2-3x=7 $$ To complete the square look at this $$ x^2+bx=c\\ (x+\frac{b}{2})^2=c+\frac{b^2}{4} $$ So in your equation what is b? Then you'll know what to add to complete the square: \(\large \frac{b^2}{4}\). Does this make sense?

OpenStudy (anonymous):

mmm im gunna be completly honest and say no , i know i must seem so dumb but i'm really really bad at math and have no clue what im doing ..